After reading all the available information carefully, prepare a two page (double-spaced) essay and answer the following questions:
Assume that we have the following data:
C=100+0.50Y
Ip=100-20r
Mt=0.10Y
Ms=100-10r
M=80

a. Build the IS-LM function.
b. If we assume an increase in Investments by 100 units, please calculate again the IS-LM functions.
c. The intersection of IS-LM functions defines four areas. Please analyze the behavior of the markets for goods and money for each area.

Ms = 100-10r
M = 80 that's the difference.
Answer #1

a. The IS–LM functions show the relationship between interest rates and real output, in the goods and services market and the money market.
The IS curve is defined by the equation Y (equilibrium of income) = C (household consumption) + I (investments) + G (government spending) + NX (Net Exports), so

Y = 100 + 0.5Y + 100 - 20r - 80,

Y = 240 - 40r.
The LM curve is defined by the equation

Mt = Ms, so

100 - 10r = 0.1Y,

Y = 1000 - 100r.

b. If we assume an increase in Investments by 100 units, then the IS function will shift to the right.

Y = 100 + 0.5Y + 100 + 100 - 20r - 80,

Y = 440 - 40r.

c. The intersection of IS-LM functions defines four areas, and the behavior of the markets for goods and money for each area will change.

Please up vote. Thanks!